# Quantum efficiency and vertical position of quantum emitters in hBN determined by Purcell effect in hybrid metal-dielectric planar photonic structures

### v1

@@ Abstract
I Introduction
Solid-state single photon emitters (SPEs) are widely seen as pivotal elements for integrated photonic quantum computing [1, 2]. They must however fulfil stringent requirements on their optical properties. This includes a high internal quantum efficiency $\eta$, which is crucial for the scalability of potential applications considering the exponential impact on the success rate of quantum protocols.
@@ I Introduction
Here, we compare the photoluminescence of well-identified individual B-centers in two different electromagnetic environments to estimate this figure of merit. Given that the intrinsic nonradiative relaxation is insensitive to the local density of states, the modifications of the decay rate allows to identify the radiative contribution to the excited state dynamics [17, 18, 19, 20]. This comparison is based on the transfer of hBN crystals hosting SPEs from a dielectric to a metallic substrate realizing a hybrid metal-dielectric planar photonic structure, which strongly modifies the emission properties of the emitters. The observed lifetime modifications are compared to numerical calculations so to infer the quantum efficiency [7], which we find close to unity. Additionally, the decay acceleration carries information about the emitter vertical position in the flake, which we estimate for multiple emitters with a confidence interval of order of a few nanometers in most cases, as established by maximum likelihood estimation. This allows to establish that stable emitters are found relatively far from the flake surface. Finally, the lossy planar cavity formed by the hybrid metal-dielectric structure allows a count rate enhancement up to a factor 8, depending on the emitter position.
II Experimental protocol and optical characterization
@@
The principle behind the present work is depicted on figure 1. The main idea is to characterize the same SPEs in two different electromagnetic environments to extract information from the variations of measured optical properties [21, 22]. The first structure consists of hBN crystals on a SiO2 (280 nm)/Si substrate (substrate A on figure 1a). The second is formed by the hBN flake and a bottom 80 nm silver film deposited on SiO2/Si (substrate B on figure 1a). More details about the sample fabrication are provided in the Methods section. The optical properties of the resulting heterostructures depend on the hBN flake thickness $h$, as well as the vertical position $d$ of the quantum emitters within the flake. The whole protocol for comparing well-identified SPEs in the two different structures is described on figure 1b. We use HPHT-grown hBN from NIMS [23], which we exfoliate on substrate A. Five flakes of thicknesses ranging from 70 nm to 220 nm are selected and irradiated using an electron beam of acceleration voltage 3-5 kV to create SPEs [10, 11], which are subsequently characterized in photoluminescence (PL). We then transfer the flakes of interest on the Ag substrate (substrate B) using a dry transfer technique (see Methods) and characterize the same emitters again in PL.
@@
The PL measurements are performed in a confocal microscope at room temperature (see Methods). A 405-nm pulsed laser is used to excite the color centers at a power of 350 $\mu$W. Lifetime and intensity autocorrelation are measured using avalanche photodiodes with a resolution of 65 ps. Additionally, the count rate is measured under cw excitation as a function of the laser power to infer the saturation count rate of the SPEs. This set of measurements is performed on well-identified SPEs before and after transfer from the SiO2 substrate to the Ag substrate. In either case, the emitter lifetime is inferred from the fluorescence decay. The decay of two representative emitters is shown on figure 2a and b, where the blue (red) curves show the measurement before (after) transfer. An exponential fit to the data provides the fluorescence lifetime $\tau_{\mathrm{A}}$ ($\tau_{\mathrm{B}}$) on substrate A (B). In the case of the first emitter, the lifetime has decreased from $\tau_{\mathrm{A}}=$2.06 ns to $\tau_{\mathrm{B}}=$1.07 ns, indicating lifetime shortening, which is attributed to the Purcell effect. For the second emitter, the lifetime is longer after transfer, increasing from $\tau_{\mathrm{A}}=$1.88 ns to $\tau_{\mathrm{B}}=$2.44 ns, thereby demonstrating Purcell inhibition of the emitter fluorescence. The statistics for $\tau_{\mathrm{A}}$ of all emitters on the SiO2 substrate is shown on figure 2d (red bars). The distribution is narrow and centered around 1.93 ns, with a standard deviation of 0.08 ns, consistently with prior work [10]. After transfer on the Ag substrate, the lifetime of most emitters is modified, resulting in the much broader distribution of $\tau_{\mathrm{B}}$ visible on figure 2d (turquoise bars). The ratio of fluorescence lifetimes $\tau_{\mathrm{A}}/\tau_{\mathrm{B}}$ measured for each emitter is plotted on figure 2e. The bar colors identify the thickness of the host flakes. For the three thinnest flakes, the lifetime is always shortened, up to a factor 2. However, in the case of the thickest flakes, the emitter decay is in some cases prolonged on substrate B with respect to substrate A. This observation of Purcell inhibition alone demonstrates that the emitter decay is not dominated by non-radiative relaxation, and allows to infer an absolute lower bound for the internal quantum efficiency $\eta$. Indeed, $\eta=\tau_{\mathrm{nr}}/(\tau_{\mathrm{rad}}^{0}+\tau_{\mathrm{nr}})$, where $\tau_{\mathrm{rad}}^{0}$ is the radiative lifetime in the absence of Purcell modification, and $\tau_{\mathrm{nr}}$ the intrinsic non-radiative lifetime. The excited state lifetime is given by $\tau^{0}=(1/\tau_{\mathrm{rad}}^{0}+1/\tau_{\mathrm{nr}})^{-1}$. In the hypothetical limit where the emitter is placed in a photonic structure that yields perfect inhibition of the radiative decay (i.e. vanishing Purcell factor), the modified lifetime is given by $\tau^{\prime}=\tau_{\mathrm{nr}}$, which corresponds to a ratio $\tau^{0}/\tau^{\prime}=\tau^{0}/\tau_{\mathrm{nr}}=1-\eta$. For finite Purcell inhibition, we then have in general $\tau^{0}/\tau^{\prime}>\tau^{0}/\tau_{\mathrm{nr}}=1-\eta$, which yields in our case $\eta>0.24$. In the next section, we perform numerical simulations to quantify the fluorescence enhancement and inhibition in our structures, allowing to provide a more accurate estimation of $\eta$.
@@
Figure 2: (a) Fluorescence decay of a SPE on substrate A (red curve) and B (blue curve), demonstrating lifetime shortening. (b) Fluorescence decay of another SPE on substrate A (red curve) and B (blue curve), demonstrating fluorescence inhibition. (c) Saturation curve of a SPE on substrate A (orange dots) and B (light blue dots). Plain curves are fits to the data (see main text). (d) Histogram of the lifetimes measured on substrate A (red bars) and substrate B (turquoise bars), showing the modification of the distribution induced by the change of photonic structure. (e) Lifetime ratio $\tau_{\mathrm{A}}/\tau_{\mathrm{B}}$ for 34 individual emitters. The bar colors indicate the thickness of the host flake. (f) Saturation count rate $C_{\infty}$ measured for several emitters on sample A (orange bars) and sample B (green bars).
@@
We also perform saturation measurements of multiple emitters on both substrates, by recording the count rate as a function of the laser power in cw regime. Figure 2c provides an example of such measurement, for the same emitter, either on substrate A (orange dots) or B (light blue dots). In this particular case, the change of substrate yielded a clear enhancement of the collection efficiency. The data are fitted using the power dependence of the count rate from a two-level system, $C(P)=C_{\mathrm{\infty}}/(1+P_{\mathrm{sat}}/P)$, allowing to extract the high-power asymptote $C_{\mathrm{\infty}}$. The coupling between the incident laser and a given emitter can vary from sample A to sample B due to interference effects between the incident and the reflected laser beam. However, this would only affect $P_{\mathrm{sat}}$ and not the asymptote $C_{\mathrm{\infty}}$, which corresponds to a maximum population of the excited state, independently of the laser power needed to reach it. Thus, $C_{\mathrm{\infty}}$ only depends on the emitter lifetime and on the collection efficiency, which also varies from sample A to B. Figure 2f presents $C_{\mathrm{\infty}}$ for several emitters, before and after transfer. The enhancement of the collection efficiency by the Ag structure varies depending on the emitter, which can be attributed to the variety of flake thicknesses and vertical positions of the measured emitters, as we confirm in the following section devoted to numerical simulations.
III Numerical simulations
To gain quantitative insight into these experimental results, we perform numerical simulations using the finite difference time domain (FDTD) method. We use a point dipole source with in-plane orientation, given that the B-centers have a single dipole oriented in the basal plane [10]. The point dipole is located at a distance $d$ from the bottom of the flake (see figure 1). To closely match our experimental protocol, for each vertical position $d$ of the emitter, we calculate the ratio $R(d)$ between the Purcell factor in both structures: $R(d)=F_{\mathrm{P}}^{B}(d)/F_{\mathrm{P}}^{A}(d)$. Experimentally, this ratio corresponds to the ratio between lifetimes $\tau_{\mathrm{A}}$ and $\tau_{\mathrm{B}}$ measured for a given emitter located at a given height $d$. Figure 3 shows $R(d)$ calculated for five flake thicknesses matching the experimentally investigated samples. In all cases, the Purcell factor diverges when $d$ vanishes, due to the proximity to the metal layer. For the three thinnest flakes, $R(d)$ has a monotonic dependence on $d$, allowing an unequivocal identification of the emitter position from the measured lifetimes $\tau_{\mathrm{A}}$ and $\tau_{\mathrm{B}}$. In the case of the two thickest flakes in turn, $R(d)$ oscillates for $d\gtrsim 50$ nm, and in some regions lies below 1, indicating a slight inhibition of the spontaneous emission. These oscillations originate from the varying coupling to the in-plane guided mode, as will be evidenced in the following section. For the 220 nm flake, we calculate the lowest value of $R(d)$ to be 0.76. This value closely matches the highest inhibition of the spontaneous emission measured for a SPE in this flake ($\tau_{\mathrm{A}}/\tau_{\mathrm{B}}=0.765\pm 0.02$, see figure 2e). Given that our simulations model a purely radiative dipole ($\eta=1$), this excellent agreement thus demonstrates that $\eta=1.00\pm 0.02$.
@@ III Numerical simulations
In the case of the 170 nm and 220 nm flakes, the Purcell factor alone does not allow a definite identification of the emitter position, since it is non-monotonic. Therefore, for the two thickest flakes, we also perform FDTD simulations to calculate the power radiated into a top objective of NA = 0.95, matching our experimental setup. Figure 4 shows the calculation results for the two crystals in configurations A and B. The collected intensity oscillates as a function of $d$ in all cases, due to a competition between in-plane and out-of-plane emission. This phenomenon is at the origin of the brightness disparities of SPEs in thick flakes observed in prior work [24]. Interestingly, the position of the nodes and antinodes is different on substrates A and B due to different boundary conditions. This leads to a position-dependent enhancement of the collection efficiency upon transfer from substrate A to B. The variation of the radiation pattern with $d$ is illustrated by the electric field profile shown on figure 4c and d, respectively at a minimum and at a maximum of the collection. In the former case, most of the light is emitted in the plane, while in the latter case, the light is directed to the top with a relatively low divergence (about 40 emission angle), which could be efficiently collected by an air objective of NA $\sim 0.65$. This shows that the metal-dielectric structure constitutes a simple yet effective way to improve the photon collection. Simulations predict collection enhancement up to a factor 8, consistent with our experimental observations. Additionally, the oscillations of the collection enhancement as a function of the emitter position provides an additional information likely to lift the remaining uncertainty about the emitter positioning as inferred only by the Purcell factor.
IV Determination of emitter positions

### v3

@@ Abstract
Introduction
Solid-state single photon emitters (SPEs) are widely seen as pivotal elements for integrated photonic quantum computing [1, 2]. They must however fulfil stringent requirements on their optical properties. This includes a high internal quantum efficiency $\eta_{i}$, which is crucial for the scalability of potential applications considering the exponential impact on the success rate of quantum protocols.
@@ Introduction
Here, we compare the photoluminescence of well-identified individual B-centers in two different electromagnetic (EM) environments to estimate this figure of merit. Given that the intrinsic nonradiative relaxation is insensitive to the local density of optical states (LDOS), the modifications of the decay rate allows to identify the radiative contribution to the excited state dynamics [17, 18, 19, 20]. This comparison is based on the transfer of hBN crystals hosting SPEs from a dielectric to a metallic substrate realizing a hybrid metal-dielectric planar photonic structure, which strongly modifies the emission properties of the emitters. The observed lifetime modifications are compared to numerical calculations so to infer the quantum efficiency [7], which we find close to unity. Additionally, the decay acceleration carries information about the emitter vertical position in the flake, which we estimate for multiple emitters with a confidence interval of order of a few nanometers in most cases, as established by maximum likelihood estimation. This allows to establish that stable emitters are found relatively far from the flake surface. Finally, the lossy planar cavity formed by the hybrid metal-dielectric structure allows a count rate enhancement up to a factor 8, depending on the emitter position.
Experimental protocol and optical characterization
The principle behind the present work is to characterize the same SPEs in two different EM environments to extract information from the variations of measured optical properties [21, 22]. In the absence of any photonic structure (i.e. in an homogeneous medium), a quantum emitter can undergo both radiative and non-radiative decay processes. We note $\tau_{\mathrm{rad}}$ the radiative lifetime, and $\Gamma_{\mathrm{rad}}=1/\tau_{\mathrm{rad}}$. The decay time associated with all intrinsic non-radiative processes, supposed independent of the LDOS, is noted $\tau_{\mathrm{nr}}$ (with $\Gamma_{\mathrm{nr}}=1/\tau_{\mathrm{nr}}$). In the case of color centers in wide-gap materials, the physical origin of such processes can include intersystem crossing to non-emissive states, as well as coupling to surrounding crystal defects and to interface- or surface-induced processes [23, 24]. The total decay rate is then $\Gamma=\Gamma_{\mathrm{rad}}+\Gamma_{\mathrm{nr}}$, and sets the excited state lifetime $\tau=1/\Gamma$. We then term $\eta_{i}$ the intrinsic quantum efficiency, defined as $\eta_{i}=\Gamma_{\mathrm{rad}}/(\Gamma_{\mathrm{rad}}+\Gamma_{\mathrm{nr}}$). This important figure of merit cannot be extracted from the fluorescence decay alone, since it requires to disentangle the two types of process. To this end, we follow the standard approach that consists of modifying the LDOS in a controlled fashion [7, 17, 18, 19, 20, 23, 24]. The subsequent changes in the emitter photophysical properties yield informations about its internal dynamics as well as its coupling to the EM field, given that only the processes associated with $\tau_{\mathrm{rad}}$ are affected, while the intrinsic non-radiative channels remain unchanged. Indeed, $\Gamma_{\mathrm{nr}}$ remains constant, while $\Gamma_{\mathrm{rad}}$ is modified to $\Gamma_{P}=\Gamma_{\mathrm{rad}}\rho(\mathbf{r})/\rho_{0}$, where $\rho_{0}$ is the LDOS in the infinite, homogeneous medium and $\rho(\mathbf{r})$ is the LDOS in the photonic structure, at a position $\mathbf{r}$. Their ratio is the Purcell factor $F_{P}(\mathbf{r})=\rho(\mathbf{r})/\rho_{0}$. We note that the new EM environment can induce additional non-radiative processes, for instance in the case of plasmonic or other metallic structures where losses can occur in the metal. These losses are included in $\Gamma_{P}$. They only act on radiative transitions and not on LDOS-independent non-radiative processes, and thus can be treated as photon losses, with no impact on the method. The modified total decay rate reads $\Gamma^{\prime}=1/\tau^{\prime}=\Gamma_{P}+\Gamma_{\mathrm{nr}}$. The intrinsic quantum efficiency can then be seen as a scaling factor that tunes the impact of a change in the LDOS on the lifetime [7], according to $\Gamma^{\prime}=\Gamma\left(1+\eta_{i}(\rho(\mathbf{r})/\rho_{0}-1)\right)$.
@@
Experimentally, we perform such comparison by successively integrating the same SPEs in two different structures, as depicted on figure 1. The first structure consists of hBN crystals on a SiO2 (280 nm)/Si substrate (substrate A on figure 1a). The second is formed by the hBN flake and a bottom 80 nm silver film deposited on SiO2/Si (substrate B on figure 1a). More details about the sample fabrication are provided in the Methods section. The optical properties of the resulting heterostructures depend on the hBN flake thickness $h$, as well as the vertical position $d$ of the quantum emitters within the flake. The whole protocol for comparing well-identified SPEs in the two different structures is described on figure 1b. We use undoped, high-purity HPHT-grown hBN from NIMS [25], which we exfoliate on substrate A. The chosen flakes exhibit a bright free-exciton peak in cathodoluminescence. This is a signature of low defect density that attributes a high quantum efficiency to the exciton, owing to limited non-radiative recombination processes [26]. We also select rather thick flakes ($h>70$ nm), which may mitigate the potential surface-induced non-radiative pathways. Five flakes of thicknesses ranging from 70 nm to 220 nm are selected and irradiated using an electron beam of acceleration voltage 3-5 kV to create SPEs [10, 11], which are subsequently characterized in photoluminescence (PL). We then transfer the flakes of interest on the Ag substrate (substrate B) using a dry transfer technique (see Methods) and characterize the sample in PL a second time. Confocal mapping allows identification of the exact same emitters before and after transfer (see Supporting Information).
@@
Figure 2: (a) Fluorescence decay of a SPE on substrate A (red curve) and B (blue curve), demonstrating lifetime shortening. The dashed black lines are exponential fits to the data. (b) Fluorescence decay of another SPE on substrate A (red curve) and B (blue curve), demonstrating fluorescence inhibition. The dashed black lines are exponential fits to the data. (c) Saturation curve of a SPE on substrate A (orange dots) and B (light blue dots). Plain curves are fits to the data (see main text). (d) Histogram of the lifetimes measured on substrate A (red bars) and substrate B (turquoise bars), showing the modification of the distribution induced by the change of photonic structure. (e) Lifetime ratio $\tau_{\mathrm{A}}/\tau_{\mathrm{B}}$ for 34 individual emitters. The bar colors indicate the thickness of the host flake. (f) Saturation count rate $C_{\infty}$ measured for several emitters on sample A (orange bars) and sample B (green bars).
The PL measurements are performed in a confocal microscope at room temperature (see Methods). A 405-nm pulsed laser is used to excite the color centers at a power of 350 $\mu$W. Lifetime and intensity autocorrelation are measured using avalanche photodiodes with a resolution of 65 ps. Additionally, the count rate is measured under cw excitation as a function of the laser power to infer the saturation count rate of the SPEs. This set of measurements is performed on well-identified SPEs before and after transfer from the SiO2 substrate to the Ag substrate. In either case, the emitter lifetime is inferred from the fluorescence decay. The decay of two representative emitters is shown on figure 2a and b, where the blue (red) curves show the measurement before (after) transfer. An exponential fit to the data (shown in dashed black lines) provides the fluorescence lifetime $\tau_{\mathrm{A}}$ ($\tau_{\mathrm{B}}$) on substrate A (B) (see Supporting Information). In the case of the first emitter, the lifetime has decreased from $\tau_{\mathrm{A}}=$2.06 ns to $\tau_{\mathrm{B}}=$1.07 ns, indicating lifetime shortening, which is attributed to the Purcell effect. For the second emitter, the lifetime is longer after transfer, increasing from $\tau_{\mathrm{A}}=$1.88 ns to $\tau_{\mathrm{B}}=$2.44 ns, thereby demonstrating Purcell inhibition of the emitter fluorescence. The statistics for $\tau_{\mathrm{A}}$ of all emitters on the SiO2 substrate is shown on figure 2d (red bars). The distribution is narrow and centered around 1.93 ns, with a standard deviation of 0.08 ns, consistently with prior work [10]. After transfer on the Ag substrate, the lifetime of most emitters is modified, resulting in the much broader distribution of $\tau_{\mathrm{B}}$ visible on figure 2d (turquoise bars). The ratio of fluorescence lifetimes $\tau_{\mathrm{A}}/\tau_{\mathrm{B}}$ measured for each emitter is plotted on figure 2e. The bar colors identify the thickness of the host flakes. For the three thinnest flakes, the lifetime is always shortened, up to a factor 2. However, in the case of the thickest flakes, the emitter decay is in some cases prolonged on substrate B with respect to substrate A. This observation of Purcell inhibition alone demonstrates that the emitter decay is not dominated by non-radiative relaxation, and allows to infer an absolute lower bound for the internal quantum efficiency $\eta_{i}$ for the emitters exhibiting fluorescence inhibition. Indeed, $\eta_{i}=\tau_{\mathrm{nr}}/(\tau_{\mathrm{rad}}+\tau_{\mathrm{nr}})$, where $\tau_{\mathrm{rad}}$ is the radiative lifetime in the absence of Purcell modification, and $\tau_{\mathrm{nr}}$ the intrinsic non-radiative lifetime. The excited state lifetime is given by $\tau=(1/\tau_{\mathrm{rad}}+1/\tau_{\mathrm{nr}})^{-1}$. In the hypothetical limit where the emitter is placed in a photonic structure that yields perfect inhibition of the radiative decay (i.e. vanishing Purcell factor), the modified lifetime is given by $\tau^{\prime}=\tau_{\mathrm{nr}}$, which corresponds to a ratio $\tau/\tau^{\prime}=\tau/\tau_{\mathrm{nr}}=1-\eta_{i}$. For finite Purcell inhibition, we then have in general $\tau/\tau^{\prime}>\tau/\tau_{\mathrm{nr}}=1-\eta_{i}$, which yields in our case $\eta_{i}>0.24$. In the next section, we perform numerical simulations to quantify the fluorescence enhancement and inhibition in our structures, allowing to provide a more accurate estimation of $\eta_{i}$.
@@
We also perform saturation measurements of multiple emitters on both substrates, by recording the count rate as a function of the laser power in cw regime. Figure 2c provides an example of such measurement, for the same emitter, either on substrate A (orange dots) or B (light blue dots). In this particular case, the change of substrate yielded a clear enhancement of the collection efficiency. The data are fitted using the power dependence of the count rate from a two-level system, $C(P)=C_{\mathrm{\infty}}/(1+P_{\mathrm{sat}}/P)$, allowing to extract the high-power asymptote $C_{\mathrm{\infty}}$. The saturation power $P_{\mathrm{sat}}$ is modified from sample A to sample B by the Purcell effect as well as by the modified propagation of the laser light due to the photonic structure, which affects the laser intensity at the emitter position due to interference effects between the incident and the reflected laser beam. The asymptote $C_{\mathrm{\infty}}$ also varies from sample A to B due to its dependence on both the radiative lifetime and the collection efficiency $\eta_{\mathrm{coll}}$ as $C_{\infty}\propto\eta_{\mathrm{coll}}/\tau_{\mathrm{rad}}$. However, since it corresponds to a maximum population of the excited state, it is independent of the laser power needed to reach it, and thus of the laser-emitter coupling, allowing to simulate the change of $C_{\infty}$ without having to account for the modification of $P_{\mathrm{sat}}$. Figure 2f presents $C_{\mathrm{\infty}}$ for several emitters, before and after transfer. The enhancement of the collection efficiency by the Ag structure varies depending on the emitter, which can be attributed to the variety of flake thicknesses and vertical positions of the measured emitters, as we confirm in the following section devoted to numerical simulations.
@@
Numerical simulations
To gain quantitative insight into these experimental results, we perform numerical simulations using the finite difference time domain (FDTD) method. We use a point dipole source with in-plane orientation, given that the B-centers have a single dipole oriented in the basal plane [10]. The point dipole is located at a distance $d$ from the bottom of the flake (see figure 1). To closely match our experimental protocol, for each vertical position $d$ of the emitter, we calculate the ratio $R(d)$ between the Purcell factor in both structures: $R(d)=F_{\mathrm{P}}^{B}(d)/F_{\mathrm{P}}^{A}(d)$. Experimentally, this ratio corresponds to the ratio between lifetimes $\tau_{\mathrm{A}}$ and $\tau_{\mathrm{B}}$ measured for a given emitter located at a given height $d$. Figure 3 shows $R(d)$ calculated for five flake thicknesses matching the experimentally investigated samples. In all cases, the Purcell factor diverges when $d$ vanishes, due to the proximity to the metal layer. For the three thinnest flakes, $R(d)$ has a monotonic dependence on $d$, allowing an unequivocal identification of the emitter position from the measured lifetimes $\tau_{\mathrm{A}}$ and $\tau_{\mathrm{B}}$. In the case of the two thickest flakes in turn, $R(d)$ oscillates for $d\gtrsim 50$ nm, and in some regions lies below 1, indicating a slight inhibition of the spontaneous emission. These oscillations originate from the varying coupling to the in-plane guided mode, as will be evidenced in the following.
For the 220 nm flake, we calculate the lowest value of $R(d)$ to be 0.76. This value closely matches the highest inhibition of the spontaneous emission measured for a SPE in this flake ($\tau_{\mathrm{A}}/\tau_{\mathrm{B}}=0.765\pm 0.02$, see figure 2e). Given that our simulations model a purely radiative dipole ($\eta_{i}=1$), this excellent agreement shows that the change of lifetime is fully accounted for by the modification of the spontaneous emission rate by the photonic structure and thus demonstrates that $\eta_{i}=1.00\pm 0.02$. To estimate the statistics of $\eta_{i}$ for the whole ensemble of emitters, we recall that in sample A (out of the cavity), the lifetime distribution is $\tau=1.93\pm 0.08$ ns. Since the statistical spread of the lifetime distribution can originate from both radiative and non-radiative processes [23], we can estimate a higher bound of the spread in $\eta_{i}$ by considering the "worst-case" scenario where the spread in $\tau$ entirely originates from a spread in $\tau_{\mathrm{nr}}$. In the limit of high quantum efficiency, the standard deviation of $\eta_{i}$ simply writes $\Delta\eta_{i}\approx\Delta\Gamma_{\mathrm{nr}}/\Gamma$, which provides $\eta_{i}=0.96\pm 0.04$ as a conservative estimation of the QE variability for the ensemble of emitters characterized in this work. Based on this argument, in the following we assume $\eta_{i}\approx 1$ for all emitters.
@@
In the case of the 170 nm and 220 nm flakes, the Purcell factor alone does not allow a definite identification of the emitter position, since it is non-monotonic. Therefore, for the two thickest flakes, we also perform FDTD simulations to calculate the power radiated into a top objective of NA = 0.95, matching our experimental setup. The simulations account for the Purcell effect, the modification of the radiation pattern, as well as losses in the metallic layers (see Methods). Therefore, the simulated collected power is proportional to the count rate collected by a two-level emitter driven at saturation. Figure 4 shows the calculation results for the two crystals in configurations A and B. The collected intensity oscillates as a function of $d$ in all cases, due to a competition between in-plane and out-of-plane emission. This phenomenon is at the origin of the brightness disparities of SPEs in thick flakes observed in prior work [27], and has already been taken benefit of in waveguide structures [15]. Interestingly, the position of the nodes and antinodes is different on substrates A and B due to different boundary conditions. This leads to a position-dependent enhancement of the collection efficiency upon transfer from substrate A to B. The variation of the radiation pattern with $d$ is illustrated by the electric field profile shown on figure 4c and d, respectively at a minimum and at a maximum of the collection. In the former case, most of the light is emitted in the plane, while in the latter case, the light is directed to the top with a relatively low divergence (about 40 emission angle), which could be efficiently collected by an air objective of NA $\sim 0.65$. This shows that the metal-dielectric structure constitutes a simple yet effective way to improve the photon collection. Simulations predict collection enhancement up to a factor 8, consistent with our experimental observations. Additionally, the oscillations of the collection enhancement as a function of the emitter position provides an additional information likely to lift the remaining uncertainty about the emitter positioning as inferred only by the Purcell factor.
Determination of emitter positions